OpenStudy (anonymous):

anyone can? Find the Taylor series expansion of a function of f(z)=z exp(2z) about z = -1 and find the region of convergence.,

OpenStudy (anonymous):

you help me? @Jemurray3

OpenStudy (anonymous):

Let u = x+1 -> x = u-1 Substitute that in first, and what do you get?

OpenStudy (anonymous):

why u = x+1 ??

OpenStudy (anonymous):

Because then you expand what you have left around u = 0

OpenStudy (anonymous):

substitute to z?

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

Oh, yeah. Z instead of x, whatever. u = z+1

OpenStudy (anonymous):

ah ok..., then f(u-1)=(u-1) exp(2(u-1)) ?

OpenStudy (anonymous):

Yep. Now just expand that in a power series around u = 0. I'll get you started. \[f(u) = e^{-2}\cdot \left( (u-1) \cdot e^{2u}\right) \] \[ =e^{-2} \cdot(u-1) \cdot \sum \frac{(2u)^n}{n!} \]

OpenStudy (anonymous):

i dont know how to expand

OpenStudy (anonymous):

Why are you having him make the substitution? That makes the problem no easier and weakens the intuition of what is happening.

OpenStudy (anonymous):

have you idea @dfyodor ??

OpenStudy (anonymous):

Well, we might as well take it from jemurray3's last equation. What is thee next part of the problem?

OpenStudy (anonymous):

the next is find the region of convergence...,

OpenStudy (anonymous):

Right, what does that mean?

OpenStudy (anonymous):

As in, the definition with equations.

OpenStudy (anonymous):

For a power series \[\sum c_n z^n\] converges if and only if \[|z|<R\], and R is called the radius of convergeance.

OpenStudy (anonymous):

We know the root test, that is, that \[\sum c_n\] converges if and only if \[\sqrt{c_n}<1\] as n goes to infinity. We also know (presumably) that \[\frac{a^n}{n!}\] converges to zero for any a greater than 0.

OpenStudy (anonymous):

I don't feel comfortable giving much more at this point, as I think that should be plenty enough to work out the solution to the problem.

OpenStudy (anonymous):

then \the region of convergence |z-1| < 1 ??

OpenStudy (anonymous):

@dfyodor I made the substitution because I'd rather that than carry around a whole bunch of (z+1)'s instead. It doesn't weaken anything. It certainly is algebraically easier to expand around zero.